### v1

@@ V Conclusion
The present micro-ARPES study, in collaboration with first-principles calculations, for three different crystal orientations of bulk RuO2 crystal has revealed the absence of altermagnetic band splitting. Besides the bulk-band degeneracy, we observed a topological surface states near associated with the bulk Dirac nodal lines irrespective of the crystal orientation. Intriguingly, its characteristics, i.e. the flat vs dispersive nature, was found to be strongly crystal-orientation dependent. The present results imply important roles of the importance of topological surface and interface states to account for the exotic physical and chemical properties of RuO2.

### v2

@@ V Conclusion
The present micro-ARPES study for three different crystal orientations, in collaboration with first-principles calculations, provides compelling spectrscopic evidence that bulk RuO2 does not exhbit altermagnetic band splitting, in agreement with recent experimental reports. Besides the bulk-band degeneracy, we observed a topological surface states near associated with the bulk Dirac nodal lines irrespective of the crystal orientation. Intriguingly, its characteristics, i.e. the flat vs dispersive nature, was found to be strongly crystal-orientation dependent, and is well captured by our slab calculations. Analysis of topological invariants, Zak phase and Wilson loop for the Dirac nodal lines suggests its non-trivial topology. The present results imply important roles of the topological surface and interface states to account for the exotic physical and chemical properties of RuO2.